# FedCCRL: Federated Domain Generalization with Cross-Client Representation Learning

### v1

@@ Abstract
Domain Generalization (DG) aims to train models that can effectively generalize to unseen domains. However, in the context of Federated Learning (FL), where clients collaboratively train a model without directly sharing their data, most existing DG algorithms are not directly applicable to the FL setting due to privacy constraints, as well as the limited data quantity and domain diversity at each client. To tackle these challenges, we propose FedCCRL, a novel federated domain generalization method that significantly improves the model’s ability to generalize to unseen domains without compromising privacy or incurring excessive computational and communication costs. Specifically, we adapt MixStyle to the federated setting to transfer domain-specific features, while AugMix is employed to perturb domain-invariant features. Furthermore, we leverage supervised contrastive loss for representation alignment and utilize Jensen-Shannon divergence to ensure consistent predictions between original and augmented samples. Extensive experimental results demonstrate that FedCCRL achieves the state-of-the-art performance on the PACS, OfficeHome, and miniDomainNet datasets across varying numbers of clients. Code is available at https://github.com/SanphouWang/FedCCRL.
@@ 1 Introduction
Domain Generalization (DG) [44] seeks to train models that can generalize effectively to unseen domains, a crucial task in real-world applications where the test data may come from domains that differ from the training data [26]. To achieve this, DG methods focus on two types of features: domain-invariant features and domain-specific features. Domain-invariant features (e.g., objects’ shape in image classification) are those that remain consistent across different domains and are essential for enabling a model to generalize to unseen domains. On the other hand, domain-specific features (e.g., color distribution of images in Figure 1) are unique to each domain and hinder generalization if overemphasized by the model.
@@
Figure 1: Illustration of the typical setting in FL. Each client contains data from a unique domain, and the test domain (Photo) differs from all domains present on the clients.
To enhance models’ generalization ability, representation learning plays a vital role in existing DG methods [1, 25, 23, 17]. A key component of this is representation alignment, which focuses on aligning feature representations of samples across domains to ensure that domain-invariant features are effectively learned. This alignment helps models capture patterns that are shared across domains while minimizing reliance on domain-specific features. In addition, feature augmentation [43, 37, 11, 41] is another powerful strategy for DG. Methods like MixStyle [43] generate new samples by mixing instance-level statistics from different domains. This approach exposes the model to a broader range of examples, thereby enhancing its ability to capture domain-invariant features while mitigating reliance on domain-specific characteristics [22], such as lighting conditions or writing styles.
We refer to these methods as centralized DG methods, as they typically assume that all data is centrally stored and accessible. However, in real-world applications, data is often distributed across multiple clients, making direct data sharing impractical due to privacy concerns. This challenge is addressed by Federated Learning (FL) [19], where clients collaboratively train models without sharing their local datasets. In order to enhance the domain generalization ability of models in FL systems, Federated Domain Generalization (FDG) [18] has been receiving increasing attention. Figure 1 provides a typical FDG setting, where each client has access to data from a single domain and the target domain differs from those on all clients. However, in such FDG setting, the limited quantity of samples and the lack of domain diversity on each client complicates the direct application of centralized DG methods [2].
Most existing FDG methods [24, 39, 7] attempt to boost the model’s generalization capacity by aligning representation distributions and extracting domain-invariant features. However, due to the limited data quantity and domain diversity on individual clients, the performance of these methods often falls short. To mitigate these shortcomings, CCST [3] leverages AdaIN [10] for cross-client style transfer to increase domain diversity on each client. However, CCST introduces additional networks and requires the transmission of a large number of high-dimensional feature representations, resulting in significant computational overhead and transmission costs. Moreover, transmitting sample features poses a considerable privacy risk, as attackers could reconstruct highly realistic samples with these feature representations and the same decoder [21, 15].
@@
To address the challenges in FDG, we propose FedCCRL—a novel method that significantly improves model generalization in the FL setting while maintaining privacy and incurring only negligible additional computational and communication costs. Specifically, FedCCRL combines cross-client feature augmentation, representation alignment, and prediction alignment. In the feature augmentation module, we adapt MixStyle [43] to the Federated Learning (FL) setting for cross-client transfer of domain-specific features, and we apply AugMix [8] to perturb domain-invariant features. For representation alignment, we utilize supervised contrastive loss [12] to align the representations of original and augmented samples. Lastly, prediction alignment leverages Jensen-Shannon divergence [20] to ensure consistency in predictions between original and augmented samples.
@@
  • We propose FedCCRL, a novel federated domain generalization framework that integrates cross-client feature augmentation, representation alignment, and prediction alignment to improve the model’s generalization capability in federated settings.
  • We adapt MixStyle to the federated learning setting, enabling cross-client style transfer while maintaining data privacy and only introducing minimal additional computational and communication costs
  • Extensive experiments on PACS, OfficeHome, and miniDomainNet datasets demonstrate that the proposed FedCCRL method achieves state-of-the-art (SOTA) FDG performance across FL systems of varying numbers of clients.
@@ 2 Related Work
This section reviews key research areas related to DG [44], with a focus on representation alignment, feature augmentation, and Federated Domain Generalization (FDG). These aspects are critical in addressing the challenges of domain shifts and ensuring robust model performance in federated learning.
@@ 2.1 Representation Alignment
One of the key techniques in representation learning for DG is representation alignment, where feature representations from different domains are aligned to minimize domain-specific variations. For instance, DANN [5, 4, 6] employs adversarial learning to align the representation distributions across domains by using a domain classifier to ensure that the learned features are domain-invariant. Another prominent approach is CORAL [28], which aligns the second-order statistics of source and target representation distributions to reduce domain shift. Additionally, MMD-based methods [29, 33, 34] utilize kernel-based distances to align representations across domains. These methods focus on reducing domain-specific discrepancies while maintaining discriminative power, helping models generalize better to unseen domains.
2.2 Feature Augmentation
Feature augmentation [31, 32, 36] has also emerged as an effective strategy for DG by generating diverse variations of the data to improve models’ robustness. One notable method is MixStyle [43], which generates new feature samples by mixing instance-level statistics (mean and variance) from different domains, effectively expanding the feature space and enabling models to generalize better across unseen domains. In addition to MixStyle, Mixup [38, 35] creates new data by linearly interpolating between two instances and their labels, using a weight sampled from a Beta distribution. Both methods are conceptually simple and computationally efficient, yet they have shown strong performance on popular benchmarks.
@@ 2.3 Federated Domain Generalization
Most existing FDG algorithms address these challenges by introducing regularization terms to enable the model to learn domain-invariant features or to ensure that feature representations across clients are aligned to a common distribution. For example, FedADG [39] utilizes federated adversarial learning to align distributions across clients to a reference distribution. FedSR [24] adopts an L2-norm regularizer and a conditional mutual information regularizer to align feature distributions across domains. In addition, FedIIR [7] takes a different approach by implicitly learning invariant relationships through gradient alignment, helping the model generalize better to OOD data. Moreover, GA [40] introduces novel global objective and optimizes it by dynamically calibrating the aggregation weights.
Although the aforementioned FDG algorithms have improved the model’s generalization ability to some extent, they do not directly address the issue posed by limited data volume and domain diversity on each client and fail to achieve the desired performance. To this end, CCST [3] employs a cross-client style transfer approach based on AdaIN [10] to generate samples with styles from other domains on the client side. However, CCST relies on a pre-trained VGG network [27] to extract feature representations and perform sample reconstruction, while also needing to transmit these high-dimensional representations, which results in significant additional communication and computational overhead. Furthermore, transmitting these feature representations violates privacy protection principles in FL, as attackers can easily intercept these representations and reconstruct the original samples with the same decoder [21, 15]. Additionally, the use of a pre-trained VGG network also partially violates the principles of domain generalization, as the target domain might be included in the pre-training dataset.
@@ 3 Method
In this section, we begin by introducing a standard FDG setting. We then describe the feature augmentation techniques utilized in FedCCRL, followed by a detailed explanation of the feature alignment regularization and the Jensen-Shannon divergence regularization. An overview of the proposed FedCCRL framework is provided in Figure 2, with the algorithmic details outlined in Algorithm 1.
@@ 3.1 Preliminary
Let $\mathcal{X}$ and $\mathcal{Y}$ denote the input space and target space, respectively. In FDG, given $M$ source domains $S_{source}=\{S_{i}|i=1,2,\dots,M\}$, the data points in $S_{i}$ are independently and identically drawn from distribution $P_{i}(x,y)$, where $x\in\mathcal{X}$, $y\in\mathcal{Y}$. The joint distribution between each pair of domains are different, i.e., $P_{i}(x,y)\neq P_{j}(x,y)$ for $1\leq i\neq j\leq M$.
@@ 3.2 Algorithm Overview
In FedCCRL, in addition to transmitting model parameters between the clients and the server, channel-wise sample statistics are also transmitted.
On the client side, for a batch of samples $X\in\mathbb{R}^{B\times C\times H\times W}$, where $B$ is the batch size, $C$ is the number of channels, and $H$ and $W$ represent the height and width of the images respectively, we first utilize feature augmentation to generate two batches of augmented samples:
@@
where $\mathcal{M}(\cdot)=\text{AugMix}\circ\text{MixStyle}$ represents the composition of MixStyle and AugMix. The detailed augmentation processes of MixStyle and AugMix are provided in Algorithm 2 and Algorithm 3, respectively.
After applying data augmentation, we extract representations and generate predictions for each batch, which are then used to compute the overall loss. Let $\mathcal{Z}$ denote the representation space. We formalize the model $f$ as $f=g\circ h$, where $h:\mathcal{X}\rightarrow\mathcal{Z}$ represents the representation encoder and $g:\mathcal{Z}\rightarrow\hat{\mathcal{Y}}$ denotes the classifier. We extract representations and produce predictions as follows:
@@
We then employ supervised contrastive loss for representation alignment, as outlined in Equation 12, and apply Jensen-Shannon divergence for prediction alignment as specified in Equation 14, to ensure consistent predictions between the original and generated samples.
@@
3.3 Feature Augmentation
In the context of DG, we identify two types of features:
  • Domain-specific features: These are features that change across domains such as color distribution, lighting conditions, or style, which have little to do with the label of the image.
  • Domain-invariant features: These features, including contours, textures, and shapes, are directly linked to the image’s label and remain consistent across different domains.
The core idea behind the feature augmentation component in FedCCRL is to perform cross-client style transfer on domain-specific features and to perturb domain-invariant features. To achieve this, we adapt MixStyle to the FDG setting and combine it with AugMix. Specifically, MixStyle handles the cross-client style transfer, while AugMix is responsible for perturbing the domain-invariant features.
MixStyle operates by leveraging the channel-wise statistics of samples. Given a sample $x\in\mathbb{R}^{C\times H\times W}$, we compute the channel-wise mean and standard deviation as follows:
@@ 3.3 Feature Augmentation
$\sigma(x)_{c}=\sqrt{\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}(x_{c,h,w}-\mu(x)_{c})^{2}}\text{.}$ (7)
In FedCCRL, in addition to transporting model parameters to the server, clients are also required to upload sample statistics. Since attackers cannot reconstruct the original samples using only channel-wise means and standard deviations, our method effectively safeguards users’ data privacy.
Specifically, each client $i$ uploads a certain proportion of sample statistics denoted as:
@@
After collecting statistics from all clients, the server concatenates these statistics to get the pool of statistics: $F_{\text{pool}}=\{F_{i}\mid i=1,2,\ldots,K\cdot M\}$, and subsequently distributes $F_{pool}^{i}=F_{pool}\setminus F_{i}$ to each client $i$. Each client apply MixStyle to generate style transferred samples according to Algorithm 2.
Algorithm 2 MixStyle on Client $i$
@@
Output: A batch of style transferred samples $X_{mixstyle}$
@@
3:     Compute $\mu(x)$, $\sigma(x)$triangleright Equation 6, 7
@@
12:$X_{mixstyle}=\text{cat}(\hat{X})$
Notably, the computational and transmission overhead associated with sample statistics is negligible compared to the costs of model training and transmission. Besides, as shown in Figure 6, our method achieves better performance than other baselines even with small upload ratio $r$. Therefore, FedCCRL does not introduce significant computational or communication overhead.
In addition to MixStyle, we employ AugMix [8] to perturb domain-invariant features. AugMix takes the original sample $x$ and processes it through multiple augmentation chains, each consisting of a sequence of simple augmentation operations. The outputs from these chains are then mixed using an element-wise convex combination. The detailed process of AugMix is provided in Algorithm 3 in supplementary material, and we adopt the same set of operations $\mathcal{O}$ as described in [8].
In Figure 3, we present the feature augmentation effects of MixStyle and AugMix. It is evident that MixStyle transfers domain-specific features, such as color distribution, while AugMix effectively perturbs domain-invariant features, such as contours of objects.
@@
(b) MisStyle
(c) AugMix
Figure 3: Effect of feature augmentation. (b) is generated from (a) using MixStyle, and (c) is generated from (b) by AugMix.

### v3

@@ Abstract
Domain Generalization (DG) aims to train models that can effectively generalize to unseen domains. However, in the context of Federated Learning (FL), where clients collaboratively train a model without directly sharing their data, most existing DG algorithms are not directly applicable to the FL setting due to privacy constraints, as well as the limited data quantity and domain diversity at each client. To tackle these challenges, we propose FedCCRL, a novel federated domain generalization method that significantly improves the model’s generalization ability without compromising privacy and without incurring excessive computational or communication costs. Specifically, we design a lightweight cross-client feature extension module that effectively increases domain diversity on each client. Furthermore, we leverage representation and prediction dual-stage alignment to enable the model to capture domain-invariant features. Extensive experimental results demonstrate that FedCCRL achieves the state-of-the-art performance on the PACS, OfficeHome and miniDomainNet datasets across FL settings of varying numbers of clients. Code is available at https://github.com/SanphouWang/FedCCRL.
@@ 1 Introduction
Traditional machine learning algorithms typically assume that training and test data follow the same independent and identically distributed (IID) pattern. However, this assumption is difficult to maintain in real-world applications, where models often encounter Out-of-Distribution (OOD) data [26]. In such cases, models trained on specific source domains (e.g., cartoon images) tend to suffer significant performance degradation on unseen domains (e.g., sketch images). To address the challenges posed by domain shift, Domain Generalization (DG) [48], which aims to enhance the model’s ability to generalize to unseen domains, has garnered significant attention. Nevertheless, traditional DG algorithms assume that the data is centrally stored and accessible to a single model, which contrasts with real-world scenarios where data is distributed across multiple clients. To enable these clients to collaboratively train a model without exchanging their raw data, Federated Learning (FL) [19] emerges as a promising framework.
@@
Figure 1: Illustration of the typical scenario in FL. Each client contains data from a unique domain, and the test domain (Photo) differs from all domains present on the clients.
However, most existing DG methods are not directly applicable to the FL setting. Specifically, traditional DG algorithms, which often rely on data sharing between domains—violating the core privacy principles of FL [21, 16]. Furthermore, as shown in Figure 1, the limited quantity of samples and the lack of domain diversity on each client further exacerbate the difficulties in achieving satisfactory performance using existing centralized DG methods [14, 45]. In order to enhance models’ domain generalization ability in FL systems, Federated Domain Generalization (FDG) [18] is emerging as an important research direction within FL.
Most current FDG methods aim to enhance the model’s generalization capability by learning domain-invariant features (i.e., label-relevant features that are consistent across domains), typically through federated adversarial learning [42, 25, 36, 20] or federated representation alignment [22, 39, 43]. Although both approaches effectively assist models in capturing domain-invariant features, they each have significant limitations. Specifically, federated adversarial learning incurs substantial computational overhead due to the utilization of discriminator. Additionally, it complicates the training process because of the risk of model collapse [1]. While federated representation alignment-based methods reduce computational overhead, they often fail to achieve satisfactory performance due to limited data quantity and domain diversity on each client, as well as the restriction against direct data sharing for privacy reasons.
Another major approach for FDG is federated style transfer [38, 40], which utilizes FL-compatible style transfer techniques, such as AdaIN [10] and CycleGAN [49], to enhance domain diversity on each client. However, most of these methods impose significant computational and transmission overheads, as they involve not only additional models for feature extraction and sample reconstruction but also the transmission of high-dimensional feature embeddings [24, 3]. Moreover, these algorithms focus exclusively on transferring domain-specific features (i.e., features that vary across domains and are minimally related to the label) and ignore the augmentation for domain-invariant features, which limits the diversity of generated samples.
@@
Inspired by the success of federated representation alignment and federated style transfer in enhancing model’s generalization ability in the FL setting, we propose FedCCRL, a novel method that significantly improves the model’s generalization capability in the FL setting while maintaining privacy and incurring only negligible additional computational and communication costs. To mitigate the adverse effects of limited data quantity and domain diversity on federated representation alignment, FedCCRL employs a lightweight cross-client feature extension module, designed to transfer domain-specific features and perturb domain-invariant features. In addition, FedCCRL applies representation and prediction dual-stage alignment to ensure the model to capture domain-invariant features.
@@
  • We propose FedCCRL, a lightweight FDG framework that effectively enhances domain diversity on each client and enables the model to capture domain-invariant features, without incurring substantial computational or transmission costs.
  • Extensive experiments on PACS, OfficeHome and miniDomainNet datasets demonstrate that the proposed FedCCRL method achieves the state-of-the-art performance across FL systems of varying numbers of clients.
@@ 2.1 Representation Alignment
A fundamental technique in representation learning-based DG methods is representation alignment, where feature representations from different domains are aligned to minimize domain-specific variations. For instance, DANN [5, 4, 6] employs adversarial learning to align the representation distributions across domains by using a domain classifier to ensure that the learned features are domain-invariant. Another prominent approach is CORAL [29], which aligns the second-order statistics of source and target representation distributions to reduce domain shift. Additionally, MMD-based methods [30, 34, 35] utilize kernel-based distances to align representations across domains.
These methods focus on reducing domain-specific discrepancies while maintaining discriminative power, helping models generalize better to unseen domains. However, most of them require the access to all of the training domains, which violates privacy principles in FL as direct data transmission between clients or to a central server is strictly prohibited.
2.2 Style Transfer
Style transfer-based DG methods [32, 33, 37] enhance domain diversity to improve model robustness across unseen domains and can be divided into two main categories.
The first category leverages generative models to synthesize data with diverse styles [27, 23], thereby reducing the model’s reliance on domain-specific features. Despite their effectiveness, these generative approaches entail substantial computational costs and are prone to issues like model collapse during adversarial training, which can hinder stable convergence.
The second category includes augmentation-based techniques, such as MixStyle [47] and Mixup [41], which diversify training data without explicit generation. MixStyle interpolates style representations within a batch, promoting domain-invariant feature learning, while Mixup creates interpolated samples that encourage generalization across domain boundaries. These approaches are more computationally efficient and avoid the instability often associated with generative models, making them advantageous for scalable DG applications.
@@ 2.3 Federated Domain Generalization
Most existing FDG algorithms rely on domain-invariant representation learning through federated adversarial learning or federated representation alignment. For instance, FedADG [42] employs a global discriminator to enable the model to extract universal feature representations across clients while holding privacy principles. In addition, FedSR [22] adopts an L2-norm regularizer and a conditional mutual information regularizer to align representation distributions across clients. However, federated adversarial learning entails significant computational overhead and the risk of model collapse. Federated representation alignment, on the other hand, demonstrates limited effectiveness in large-scale federated learning contexts due to insufficient domain diversity on individual clients.
While these FDG algorithms have contributed to enhancing the model’s generalization ability to a certain extent, they remain insufficient in effectively addressing the constraints of limited data volume and domain diversity on each client, which impedes optimal performance. To this end, CCST [3] employs a cross-client style transfer approach based on AdaIN [10] to generate samples with styles from other domains on the client side. However, CCST relies on a pre-trained VGG network [28] to extract feature representations and perform sample reconstruction, while also needing to transmit these high-dimensional representations, which results in significant additional communication and computational overhead. Furthermore, transmitting these feature representations violates privacy protection principles in FL, as attackers can easily intercept these representations and reconstruct the original samples with the same decoder [21, 16]. Additionally, the use of a pre-trained VGG network also partially violates the principles of domain generalization, as the target domain might be included in the pre-training dataset.
Apart from the methods discussed above, other FDG approaches take different strategies. For instance, FedIIR [7] implicitly learns invariant relationships across domains by aligning gradients, which allows the model to generalize to OOD data. Besides, GA [44] aids FDG by dynamically calibrating aggregation weights, thus minimizing generalization gaps across clients. However, these methods have not demonstrated significant advantages over other FDG approaches in experimental performance.
@@ 3.1 Preliminary
Let $\mathcal{X}$ and $\mathcal{Y}$ denote the input space and target space respectively. In FDG, given $M$ source domains $S_{source}=\{S_{i}|i=1,2,\dots,M\}$, the data points in $S_{i}$ are independently and identically drawn from distribution $P_{i}(x,y)$, where $x\in\mathcal{X}$, $y\in\mathcal{Y}$. The joint distribution between each pair of domains are different, i.e., $P_{i}(x,y)\neq P_{j}(x,y)$ for $1\leq i\neq j\leq M$.
@@ 3.2 Algorithm Overview
We provide the overview of the proposed FedCCRL framework in Figure 2, with the algorithmic details outlined in Algorithm 1.
On the client side, for a batch of samples $X\in\mathbb{R}^{B\times C\times H\times W}$, where $B$ is the batch size, $C$ is the number of channels, and $H$ and $W$ represent the height and width of the images respectively, we first apply cross-client feature extension to generate two batches of augmented samples:
@@
where $\mathcal{M}(\cdot)=\text{CCDT}\circ\text{DIFP}$ represents the composition of CCDT and DIFP in the cross-client feature extension module. The detailed procedures for these components are provided in Algorithm 2 and Algorithm 3 respectively.
After cross-client feature extension, we extract representations and generate predictions for each batch, which are then used to compute the overall loss. Let $\mathcal{Z}$ denote the representation space. We formalize the model $f$ as $f=g\circ h$, where $h:\mathcal{X}\rightarrow\mathcal{Z}$ represents the representation encoder and $g:\mathcal{Z}\rightarrow\hat{\mathcal{Y}}$ denotes the classifier. We extract representations and produce predictions as follows:
@@
Subsequently, we utilize supervised contrastive loss for representation alignment, as defined in Equation 12, and employ Jensen-Shannon divergence for prediction alignment, as specified in Equation 14, to ensure consistent predictions between the original and generated samples.
@@
3.3 Lightweight Cross-Client Feature Extension
While multiple studies [2, 15] have shown that combining domain-specific and domain-invariant feature augmentations can create a more adaptable feature space, most data augmentation-based FDG methods transfer only domain-specific features and often introduce significant computational and transmission overhead.
In contrast, FedCCRL introduces a lightweight cross-client feature extension module comprising two components: Cross-Client Domain Transfer (CCDT) and Domain-Invariant Feature Perturbation (DIFP). This cross-client feature extension module effectively transfer samples’ domain-specific features across clients and perturbs domain-invariant features without incurring excessive computational and transmission costs.
CCDT, inspired by MixStyle [47], leverages sample statistics to transfer domain-specific features. Specifically, in FedCCRL, clients are required not only to transmit model parameters to the server but also to upload the channel-wise means and standard deviations of samples:
@@ 3.3 Lightweight Cross-Client Feature Extension
$\sigma(x)_{c}=\sqrt{\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}(x_{c,h,w}-\mu(x)_{c})^{2}}\text{,}$ (7)
where $x\in\mathbb{R}^{C\times H\times W}$ is a given sample.
Each client $i$ uploads a certain proportion of sample statistics denoted as:
@@
After collecting statistics from all clients, the server concatenates these statistics to get the pool of statistics: $F_{\text{pool}}=\{F_{i}\mid i=1,2,\ldots,K\cdot M\}$, and subsequently distributes $F_{pool}^{i}=F_{pool}\setminus F_{i}$ to each client $i$. Each client apply CCDT to generate style transferred samples according to Algorithm 2.
Algorithm 2 CCDT on Client $i$
@@
Output: A batch of style transferred samples $X_{CCDT}$
@@
3:     Compute $\mu(x)$, $\sigma(x)$triangleright Eq. 6, 7
@@
12:$X_{CCDT}=\text{cat}(\hat{X})$
In DIFP, we employ AugMix [8] to perturb domain-invariant features. AugMix takes the original sample $x$ and processes it through multiple augmentation chains, each consisting of a sequence of simple augmentation operations. The outputs from these chains are then mixed using an element-wise convex combination. The detailed process of AugMix is provided in Algorithm 3 in supplementary material, and we adopt the same set of operations $\mathcal{O}$ as described in [8].
In Figure 3, we present the effects of CCDT and DIFP. It is evident that CCDT transfers domain-specific features, such as color distribution, while DIFP effectively perturbs domain-invariant features, such as contours of objects.
@@
(b) CCDT
(c) DIFP
Figure 3: Effect of feature augmentation. (b) is generated from (a) by CCDT, and (c) is generated from (b) by DIFP.